Monday, March 27, 2006

Duplicate IP for Dorm Students

About two weeks ago this issue surfaced. We've been working to resolve it. For anyone encountering the problem, we have had some success with using the IPCONFIG /RELEASE and then the IPCONFIG /RENEW command from the command prompt.

Voice Mail Stablized

We believe the voice-mail system is stabilized at this time. If you encounter further issues please let us know through one of the helpdesk channels ( - the preferred way).

Friday, March 24, 2006

Voice Mail

We are experiencing intermittent problems with the Voice Mail System. A service call has been placed.

Friday, March 17, 2006

TItanNet Group and Home Directory Issue

This morning we encountered an issue reaching group and home directories. Thanks to those who reported the matter. Access has been restored.

This has been a problem for a little while and we are continuing to research the matter to bring it to conclusion.

Thursday, March 09, 2006

Listserver Issue

The listserver is presently not delivering messages due to a technical issue (disk space related). Staff have it back up but and the queues are now processing but there are a number of messages that have yet to process through.

Outer Drive Internet Down

A scheduled upgrade of the WIFI bridge is taking place today and tomorrow. Access at the new and improved rate should be available after 3:00 pm on 3/10/06.

Friday, March 03, 2006

Campus Wide Internet Access Restored

Approximately 4:30 pm all power in Shiple brought failed and brought down the wireless bridge to our ISP. This was related to work on the fire alarm systems. Power and internet access was restored about 7:00 pm.